A long-distance runner has an average speed of 2 m/s during a race. How far does the runner travel in 25 min?
Answer #1

distance=speed*time

distance=2*25*60=3000m or 3 km
Answer #2

Distance = Average Speed * Time

= 2 * (25 * 60)

= 3000 m
Answer #3

s=v*t=2*25*60=3000 m= 3 Km